OpenStudy (anonymous):

I have a question for Algebra II. I have a question for Algebra II. @Mathematics

OpenStudy (anonymous):

Here is the attatchment

OpenStudy (anonymous):

I take an online course in Math because I am homeschooled.

OpenStudy (saifoo.khan):

that's cool. :p

OpenStudy (saifoo.khan):

\[-0.5 x+y=-2\]

OpenStudy (asnaseer):

standard form of a straight line is y = mx + c you are given m=-1/2, so you know y = -x/2 +c so now you just need to find "c" you are also told that the line passes through the point (-2,-3) so you know that when x=-2, y=-3. this should be enough for you to solve this.

OpenStudy (anonymous):

Ok thanks guys. I have couple of other questions, that I need to know how to solve them.

OpenStudy (anonymous):

Thats slope-intercept asnaseer.. Standard form is Ax +By = C

OpenStudy (anonymous):

Yeah, and I don't know how to form the equation.

OpenStudy (anonymous):

\[\large -\frac{1}{2}x - 3y = -2\]

OpenStudy (anonymous):

I think thats right.

OpenStudy (saifoo.khan):

wait, mine isnt wrong.

OpenStudy (anonymous):

then mine is wrong lol

OpenStudy (anonymous):

If you look at the attatchment... it will not allow me to put that equation in.

OpenStudy (anonymous):

One of you is right, lol

OpenStudy (anonymous):

its probably not me lol

OpenStudy (anonymous):

Ok, so I can't put it in the system though.

OpenStudy (anonymous):

wlbrt316 What do you think?

OpenStudy (asnaseer):

using the method I outlined above, I get c=-4, so y = -x/2 - 4 which can be re-written as: x + 2y = -8

OpenStudy (anonymous):

Ok I am trying it as I speak... Lol So hold on?

OpenStudy (anonymous):

Good answer... I got 4/5 correct... can you help me out with another?

OpenStudy (anonymous):

I have another... is anyone up for it?

OpenStudy (anonymous):

Question 19 is kind of getting me stuck.